find the value of k so that the line containing the points (-4,k) and (2,8) is perpendicular to the line containing the points (8,-3) and (-6,9).
Answer by scott8148(6628)

the slope of the known line is (9-(-3))/(-6-8) or 12/(-14) or -6/7

prependicular means the solpe is the negative reciprocal or 7/6

(8-k)/(2-(-4))=7/6 __ (8-k)/6=7/6 __ 8-k=7